in RC coupled amplifier the battery voltage is 16V and collector load Rc=4kΩ. its required to set the operating point at Ic=1mA , VCE= 10V by potential divider method. if VBE= 0.2V and I1 = 10IB ẞ= 100 find various circuit values. 


1
Expert's answer
2020-09-30T12:22:58-0400

Given

"I_c=1mA,V_{ce}=10 V,R_c=4K\Omega, \beta=100,V_{BE}=0.2V"

Battery voltage "V_{BB}=16 V"

Using

"V_{CE}=V_{CC}-I_cR_c"


"10=V_{CC}-1\times 4"

"V_{cc}=14 V"

we know

"\beta=\frac{I_c}{I_b}"


"I_b=\frac{1}{100}"


=0.01mA

using

"V_{BE}=V_{BB}-I_bR_b"


"0.2=16-0.01R_b"

so "R_b=\frac{15.8}{0.01}"

="1580 K\Omega"

"I_E=I_b+I_c"

=0.01+1

=1.01mA